# Fair Wasserstein Coresets

### v3

@@ C.2 Real Datasets
Figure 6: Fairness-utility tradeoff of all methods, indicated by AUC and demographic disparity of a downstream MLP classifier across all datasets (rows) and without (left column) or with (right column) fair pre-processing Kamiran & Calders (2012) (excluding FWC , to which no fairness modification is applied after coresets have been generated). We report the averages over 10 separate train/test splits.
@@
Figure 7: Similarly to Figure 6, we report the means and standard deviations over 10 runs of the fairness-utility tradeoff of all methods, indicated by AUC and demographic disparity of a downstream MLP classifier across all datasets (rows) and without (left column) or with (right column) fair pre-processing Kamiran & Calders (2012) (excluding FWC , to which no fairness modification is applied after coresets have been generated).

### v4

@@ Appendix E Broader Impact
Figure 6: Fairness-utility tradeoff of all methods, indicated by AUC and demographic disparity of a downstream MLP classifier across all datasets (rows) and without (left column) or with (right column) fair pre-processing [34] (excluding FWC , to which no fairness modification is applied after coresets have been generated). The Pareto frontier, indicated with a dashed red-line, is computed across all models and coreset sizes. We report the averages over 10 separate train/test splits.
@@
Figure 7: Similarly to Figure 6, we report the means and standard deviations over 10 runs of the fairness-utility tradeoff of all methods, indicated by AUC and demographic disparity of a downstream MLP classifier across all datasets (rows) and without (left column) or with (right column) fair pre-processing [34] (excluding FWC , to which no fairness modification is applied after coresets have been generated).